OpenStudy (anonymous):

HELP PLEASE BEST ANSWER GETS THE MEDAL............................................................................ Which equation is a quadratic equation? y + 7 = −3x(4 − x) + 9x y = −2x + 5 y + 2x3 = (2x − 2)(x + 5) y − 9x = (x2 + 1)(5x − 1) + 3

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

y + 7 = −3x(4 − x) + 9x y=-12x+3(x^2)+9x-7 y=3x^2-3x-7 so because of x^2 it's quadratic equation